OpenStudy (anonymous):

What are the three consecutive odd integers of: x + (x+2) + (x+4) = 537 A. 177, 178, 179 B. 177, 179, 181 C. 181, 182, 183 D.181, 183, 185

zepdrix (zepdrix):

X represents the FIRST of the three odd integers. So simply solve for X, and you will have found the first odd number. The next number will be 2 MORE than X as described by the (x+2). \[\large x+(x+2)+(x+4)=537 \quad \rightarrow \quad x+x+2+x+4=537\]\[\large 3x+6=537\]Understand how I combined the x's?